what is the radius of the circle increases at the rate of 0.01 inch per second, find the rate of change of the area when the radius is 3 inches long


1
Expert's answer
2021-12-20T16:55:06-0500

We know that the area of the circle determine as "S=\u03c0r^2." Also we have "\frac{dr}{dt}=0.01" in/sec, because the radius of the circle increases at the rate of 0.01 in per second.

Find "\frac{dS}{dt}=2πr\frac{dr}{dt}."


Such as r=3 in, then


 "\frac{dS}{dt}=2π•3•0.01=0.06π" in2/sec or 0.1884 in2/sec


Answer. 0.1884 in2/sec